Candy worth $2.55 per pound is mixed with candy worth $2.40 per pound. How much of each kind must be used to have 22.5 pounds of a mixture worth $2.48 per pound?
Answer by mananth(16946)

---------price ---------------- quantity
Candy I 2.55 ---------------- x pounds
Candy II 2.40 ------ 22.5 - x pounds
Mixture 2.48 ---------------- 22.5

---------price ---------------- quantity
Candy I 2.55 ---------------- x pounds
Candy II 2.4 ------ 22.5 - x pounds
Mixture 2.48 ---------------- 22.5

2.55 x + 2.4 ( 22.5 - x ) = 2.48 * 22.5

2.55 x + 54 - 2.4 x = 55.8
2.55 x - 2.4 x = 55.8 - 54
0.15 x = 1.8
/ 0.15
x = 12 pounds $2.55 Candy I
10.5 pounds $2.40 Candy II